OpenStudy (anonymous):

Find the missing diagonal of a rhombus if the area is 92ft2 and the other diagonal is 8 feet

OpenStudy (anonymous):

23

OpenStudy (vishweshshrimali5):

d1 d2 / 2 = 92 => d1 * 8 = 92*2 => d1 = 184/8 = 23

OpenStudy (anonymous):

area=0.5*8*dia 92=4*dia dia=23